Status Report

ISS On-Orbit Status 3 Jan 2003

By SpaceRef Editor
January 3, 2003
Filed under , ,
ISS On-Orbit Status 3 Jan 2003

All ISS systems continue to function nominally, except as noted previously or below.  

CDR Ken Bowersox stowed the acoustic dosimeters used in the last few days for an internal noise survey.  

Later, FE-2/SO Don Pettit was to prepare Russian SLM (sound level meter) equipment and to take readings at 41 different locations in the Lab, Node, Airlock, FGB, Service Module (SM) and Docking Compartment (DC-1), following instructions uplinked on 1/1.  The noise measurement data were then to be transferred to the MEC (medical equipment computer).

FE-1 Nikolai Budarin performed the daily checkup of the BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) experiment for the research of plant growth and development under spaceflight conditions.

Afterwards, Nikolai spent an hour in the DC-1, cleaning the D3 air return vent grille, location of an HEPA (high-efficiency particulate air) filter.

Bowersox had a full day of ZCG (Zeolite crystal growth) payload activities, consisting of IZECS (improved Zeolite electronic control system) powerdown and subsequent power cycling, HDD (hard disk drive) replacement, furnace unloading and later activation, autoclave configuring, bubble coalescing, furnace heater activation, operation monitoring, and functionality check.  Sox was then to stow parts and tools temporarily until the end of the day. [Because the EXPRESS Rack 2 (ER2) laptop is not communicating with the RIC (rack interface controller), the ground was to command power and comm off/on throughout the day.]

The crew was asked to check the power supplies of the UOP-3 (utility outlet panel 3) and the MACE Silo II (no. 2 mass storage device for Ops LAN file server) on a time-available basis.  [The RPC (remote power controller) for Lab UOP-3 tripped on 1/1 at about 2:12pm EST. Telemetry from both the RPCM (remote power controller module) and the upstream DDCU (dc-to-dc current converter) indicated elevated currents at multiple instances prior to the trip. It is unlikely that the UOP’s 28 Volt loads (camera chargers, bar code reader battery charger, SMPA/battery charger, etc.) or the PS-28 power strip caused the problem. The two MACE II power strips will convert 500 Watts maximum, while 1500 W is required to trip the RPCM.  The problem is suspected either within the MACE II power supplies, UOP, or RPCM.  The crew was advised to continue use of these items at another UOP.]

Yesterday, Bowersox started the deferred A/L VAJ (Airlock vacuum access jumper) leak check, and later today he was scheduled to terminate it.  Afterwards, he was to address some questions concerning A/L items resulting from the PBA/PFE (portable breathing apparatus/portable fire extinguisher) audit performed on 12/31/02.

FE-1 Budarin conducted an inventory of the various DVCAM, Betacam and LIV videotapes on board at various stowage locations in the SM, DC-1 and FGB.

Ken Bowersox was scheduled to work on the ER2 ARIS (active rack isolation system), disassembling and removing the snubber cup guides.

Don Pettit transferred condensate water from the Lab condensate collection tank to CWC (contingency water container) and then took a CWC inventory.

Bowersox collected air samples with the VOA (volatile organics analyzer), while Pettit took photo documentation of the sampling, followed by closeout activities.

Budarin’s work schedule included the task to switch the SM audio subsystem (STTS), with its telephone/telegraph audio paths and comm stations, to its primary string.  [The “Voskhod-M” STTS enables telephone communications between the SM, FGB, DC-1 and USOS (US segment), and also with users on the ground via STTS antennas on the outside of the SM and VHF channels selected by an operator at the SM comm panel.  There are six comm panels in the SM with pushbuttons for accessing any of three audio channels plus an intercom channel.]

The daily routine maintenance of SOSh life support systems, incl. ASU toilet facility, was performed by Budarin, while Bowersox completed the regular daily Lab payload status checkup, and Pettit prepared the IMS (inventory management system) delta file for downlinking.

During yesterday’s OBT (on-board training) for fire emergency procedures, it was reported that the audio between Russian segment and US segment was not working for Public Loop 3. The crew was requested today to provide MCC-H with the configuration of SM Comm Panel 2 during this activity and to report if the problem has been resolved or not.

Expedition 6 is providing a wealth of digital imagery for download via OCA.  Current net rate of picture-downloading is about 236 images/day (i.e., 1.6 gigabytes/day).  Up to twice that number is stored on board for later downlink.

All crewmembers performed their daily physical exercise program.

Today’s targets for the CEO (crew earth observations program) were Tashkent, Uzbekistan (nadir pass over this capital city in the foothills of the Tien Shan Mountains.  ESC [electronic still camera] requested.  Obliques should capture regional smoke that is making the news.  The most comprehensive and long term study of global burning [based on handheld photos taken over the last 20 yr] shows that burning in high latitude Asia starts in March:  regional smoke in January is therefore noteworthy), Bangkok, Thailand (nadir pass; ESC), Bombay, India (nadir pass; ESC.  Crew was to shoot city boundaries [in one or two near-nadir views, if possible], and any smoke/air pollution drifting out to sea [obliques looking right are best]), Saharan stream patterns, Niger (looking 2 degrees right of track for ancient stream patterns in the Saharan desert floor, between ISS track and the line of the Niger River [running parallel to the track].  These major streams flowed repeatedly in the past when the Sahara was wetter [roughly coincident with the ice ages in higher latitudes]), Albuquerque, New Mexico (nadir pass; ESC), Lower Rio Grande, Texas (mapping strip from Big Bend to Brownsville requested in this nadir pass.  Crew was to shoot both sides of the border in this fastest growing part of the US). Melbourne, Australia (nadir pass; ESC), and SE Sudan stream patterns (sun glint opportunity is ideal for documenting the highly complex stream patterns in swamplands from the Nile River along the Ethiopian mountain front [Sudan-Ethiopia border] as far as Lake Turkana [~2 min].  Mapping swath half right requested.  Then the crew was to shoot the Omo River delta [at nadir] at the north end of the major Lake Turkana.  The exposed area of the delta has changed repeatedly with the rise and fall of this inland lake, which is an indicator of regional rainfall, related mainly to El Nino oscillations.  A recent image to continue the time series is requested).

SpaceRef staff editor.